Question:
Find the sum of the grometric series: 16, -4, 1, -1/4...
(A) 4415/320
(B) 4095/320
(C) 64/5
(D) 16,383/1280
Im stuck- thanks for your help!

Answer by venugopalramana(3286)

SUM OF A G.P.TO INFINITE TERMS,WHERE COMMON RATIO IS LESS THAN 1 IS GIVEN BY
S=A/(1-R)...WHERE A IS FIRST TERM AND R = C.R.
HERE A=16 AND R = -4/16=-1/4
S=16/(1+1/4)=16/(5/4)=16*4/5=64/5